OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -2 - 6t. Find the instantaneous velocity at t = 2 by finding the derivative.

OpenStudy (anonymous):

@Mertsj Please?

OpenStudy (isaiah.feynman):

Take the derivative of the position function.

OpenStudy (anonymous):

Apply t=2 to -2-6(2)?

OpenStudy (isaiah.feynman):

@kokopuffs890 no that would be wrong. You need to take the derivative first.

OpenStudy (anonymous):

ok i know the derivative for this is -6, i just dont know how to show it

OpenStudy (anonymous):

@Isaiah.Feynman

OpenStudy (isaiah.feynman):

I'm working on it. I used this \[\lim_{h \rightarrow 0} \frac{ f(t+h)-f(t) }{ h }\]. But I got 6 which is not correct yet.

OpenStudy (anonymous):

OH i see... i got it from here: http://www.wolframalpha.com/input/?i=derivative%3A+-2-6t%2C+t%3D2

OpenStudy (isaiah.feynman):

Yes the derivative is indeed -6 and you can use that formula I used to show it. That formula is called the definition of a derivative.

OpenStudy (anonymous):

So by just showing to solve that, would be correct the answer?

OpenStudy (isaiah.feynman):

Yeah, use the definition to solve it.

OpenStudy (anonymous):

ok lets see...

OpenStudy (anonymous):

what is h in the equation? -6?

OpenStudy (isaiah.feynman):

Okay, lets start afresh, discard everything I already said. Now the definition of a derivative is \[\lim_{h \rightarrow 0} \frac{ f(a+h) - f(a) }{ h }.\]. In that equation a is 2, and t=a+h. Substitute these values and see what happens.

OpenStudy (anonymous):

ok, f(2+h)-f(2) / h

OpenStudy (isaiah.feynman):

Yes and remember t=2+h.

OpenStudy (anonymous):

s(t)=2-6t \[\frac{ ds }{dt }=-6,at t=2,velocity = -6\]

OpenStudy (anonymous):

so t-2=h?

OpenStudy (anonymous):

oh wait! that formula, right, that would work too right @Isaiah.Feynman ?

OpenStudy (isaiah.feynman):

The formula I just gave would work!

OpenStudy (anonymous):

ok cool, so t-2 = h? if we substitute?

OpenStudy (isaiah.feynman):

Yeah, but you don't have to substitute though. \[\lim_{h \rightarrow 0} \frac{ -2-6(2+h)-(2-12) }{ h }\], simplifying gives us \[\lim_{h \rightarrow 0}\frac{ -2 -12-6h+2+12 }{ h }\] 2's and 12's cancel out leaving us with

OpenStudy (isaiah.feynman):

\[\lim_{h \rightarrow 0} \frac{ -6h }{ h } = -6\]

OpenStudy (anonymous):

Oh wow ok, now i get it, and this methods always works?

OpenStudy (isaiah.feynman):

Yeah, but when the function looks wingspan eyed, you can use the differentiation rules.

OpenStudy (anonymous):

right i see, ok great, thanks for the help friend :D

OpenStudy (isaiah.feynman):

No probs.